A plane took 1 hour longer to travel 560 miles on the first portion of a flight than it took to fly 480 miles on the second portion. If the speed was the same for each portion, what was the flying time for the second part of the trip?
Answer by rajagopalan(133)

Let time taken for second part be = x hours
Then time taken for First part =x+1 hours
Now From given data 560 miles in x+1 hours...Eqn 1
Now From given data 480 miles in x hours...Eqn 2
Subtracting 080 miles in (x+1)-x=1 hour
The speed of plane is 80 miles per hour
Time for second portion = 480/80 = 6 hours
OK